OpenStudy (anonymous):

Solve 2 over 7 = 6 over x

OpenStudy (anonymous):

2/7=6/x cross multiply 2x=6*7 2x=42 solve for x x=21

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

easy way just change x put at up side.. 7/2=x/6